You will combine two numbers from 0 to 9. One entry will win PHP 6000, out of 100 total entries, if 

you match the last two digits of the final score in a basketball game. Let X be the net gain from 

the purchase of one entry. What is the mean or the expected value of a person who buys a ticket? 

Illustrate by constructing a probability distribution. 



1
Expert's answer
2021-05-17T18:51:15-0400

Total number of ways in which we can choose two numbers "=^{10}C_2=45"


On entry will win PHP 6000


Probability that will win the entry "=\dfrac{45}{100}=0.45"


Probability that does not win "= \dfrac{55}{100}=0.55"


Probability distribution is-




Expected value of person who buys a ticket = "0.

45\times 6000+0.55\times 0"

"=2700"


Hence Expected amount of winning is php 2700.